a random variable x observes the following density in the interval [0, 1] f(x) = { a + bx, for 0 ≤ x ≤ 1 0, elsewhere .for what value of a and b does the median associated with the above density equal 1/2


1
Expert's answer
2021-07-20T12:32:47-0400
"f(x) = \begin{cases}
 a+bx &0\leq x\leq 1 \\
 0 &elsewhere
\end{cases}"

"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1}(a+bx)dx=[ax+\dfrac{1}{2}bx^2]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=a+\dfrac{1}{2}b=1"

"\displaystyle\int_{0}^{1/2}(a+bx)dx=[ax+\dfrac{1}{2}bx^2]\begin{matrix}
 1/2 \\
 0
\end{matrix}"

"=\dfrac{1}{2}a+\dfrac{1}{8}b=\dfrac{1}{2}"


"a+\dfrac{1}{2}b=1"

"a+\dfrac{1}{4}b=1"

"a=1, b=0"